# General Coded Computing in a Probabilistic Straggler Regime

### 08 Mar 2026 v2

@@ Abstract
Coded computing has demonstrated promising results in addressing straggler resiliency in distributed computing systems. However, most coded computing schemes are designed for exact computation, requiring the number of responding servers to exceed a certain recovery threshold. Additionally, these schemes are tailored for highly structured functions. Recently, new coded computing schemes for general computing functions, where exact computation is replaced with approximate computation, have emerged. In these schemes, the availability of additional results corresponds to more accurate estimation of computational tasks. This flexibility introduces new questions that need to be addressed. This paper addresses the practically important scenario in the context of general coded computing, where each server may become a straggler with a probability $p$, independently from others. We theoretically analyze the approximation error of two existing general coded computing schemes: Berrut Approximate Coded Computing (BACC) and Learning Theoretic Coded Computing (LeTCC). Under the probabilistic straggler configuration, we demonstrate that the average approximation error for BACC and LeTCC converge to zero with the rate of at least $\mathcal{O}(\log^{3}_{\nicefrac{{1}}{{p}}}(N)\cdot{N^{-3}})$ and $\mathcal{O}(\log^{4}_{\nicefrac{{1}}{{p}}}(N)\cdot{N^{-2}})$, respectively. This is perhaps surprising, as earlier results does not indicate a convergence when the number of stragglers scales with the total number of servers $N$. However, in this case, despite the average number of stragglers being $Np$, the independence of servers in becoming stragglers allows the approximation error to converge to zero. These theoretical results are validated through experiments on various computing functions, including deep neural networks.
@@ I Introduction
In this paper, we address this question and show that the above understanding is not correct. We theoretically analyze the approximation error in the presence of probabilistic stragglers for two general coded computing schemes of BACC and LeTCC. We demonstrate that the approximation errors of BACC and LeTCC schemes converge to zero at rates of at least $\mathcal{O}(\log^{3}_{\frac{1}{p}}(N)N^{-3})$ and $\mathcal{O}(\log^{4}_{\frac{1}{p}}(N)N^{-2})$, respectively. This result demonstrates that the independence of servers in becoming stragglers indeed contributes positively to the convergence of the error. In the end, we validate our theoretical findings through experiments on various computing functions, including deep neural networks (see Fig. 1).
@@ III Problem Formulation
Following the notation from [6], the estimator function for the coded computing scheme is denoted by $\mathbf{\hat{f}}_{\bm{\alpha},\bm{\beta}}[\mathbf{u}_{\textrm{enc}},\mathbf{u}_{\textrm{dec}},\mathcal{F}](\cdot)$. where, $\bm{\alpha}=[\alpha_{1},\dots,\alpha_{K}]^{T}$ and $\bm{\beta}=[\beta_{1},\dots,\beta_{N}]^{T}$ and $\mathcal{F}$ refers to the set of non-straggler servers.
We define a random variable $F_{N}$ that is distributed over the set of all subsets of $N$ servers. Consequently, the performance of a coded computing estimator $\mathbf{\hat{f}}(\mathbf{x}):=\mathbf{\hat{f}}_{\bm{\alpha},\bm{\beta}}[\mathbf{u}_{\textrm{enc}},\mathbf{u}_{\textrm{dec}},\mathcal{F}](\cdot)$, applied to the data points $\{\mathbf{x_{k}}\}_{k=1}^{K}$, can be evaluated using the following average approximation error:
@@ IV Main Results
$\mathcal{L}(\hat{f})\leqslant C_{l}\frac{\left(\log_{\frac{1}{p}}(qN)+\sqrt{\frac{1}{\delta}}\right)^{3}}{N^{3}}\cdot\left\lVert(f\circ u_{\textrm{enc}})^{\prime\prime}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}+\frac{2\nu^{2}}{K}\sum_{k=1}^{K}\left(f(u_{\textrm{enc}}(\alpha_{k}))-f(x_{k})\right)^{2}.$ (7)
@@
$\mathcal{L}(\hat{f}){\leqslant}C_{b}\cdot B\frac{\left(\log_{\frac{1}{p}}(qN)+\sqrt{\frac{1}{\delta}}\right)^{4}}{N^{2}}\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}.$ (8)
@@
By applying Theorems 1 and 2, in the probabilistic straggler setting, the average approximation error converges to zero with high probability. The convergence rates are at least $\mathcal{O}(\frac{\log^{3}_{\frac{1}{p}}(N)}{N^{3}})$ for LeTCC and at least $\mathcal{O}(\frac{\log^{4}_{\frac{1}{p}}(N)}{N^{2}})$ for BACC.
@@ V Proof Sketch
$\mathcal{L}(\hat{f})\leqslant\underbrace{\mathop{\mathrm{E}}_{\mathcal{F}\sim F_{N}}\left[\frac{2}{K}\sum^{K}_{k=1}\left(u_{\textrm{dec}}(\alpha_{k})-f(u_{\textrm{enc}}(\alpha_{k}))\right)^{2}\right]}_{\mathcal{L}_{\textrm{dec}}(\hat{f})}+\underbrace{\frac{2}{K}\sum^{K}_{k=1}\left(f(u_{\textrm{enc}}(\alpha_{k}))-f(\mathbf{x}_{k})\right)^{2}}_{\mathcal{L}_{\textrm{enc}}(\hat{f})}$. (9)
@@
$\mathrm{E}_{\mathcal{F}\sim F_{N}}\left[\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\right]\leqslant 2\mathrm{E}_{\mathcal{F}}\left[\left\lVert h\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\right]^{\frac{1}{2}}\cdot\mathrm{E}_{\mathcal{F}}\left[\left\lVert h^{\prime}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\right]^{\frac{1}{2}}.$
@@
$\mathrm{E}_{\mathcal{F}}\left[\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\right]\leqslant C\cdot D\mathrm{E}_{\mathcal{F}}\left[(\Delta^{\mathcal{F}}_{\textrm{max}})^{4}\right]^{\frac{1}{2}}\cdot\sqrt{D}\mathrm{E}_{\mathcal{F}}\left[(\Delta^{\mathcal{F}}_{\textrm{max}})^{2}\right]^{\frac{1}{2}}\cdot\sqrt{H_{0}H_{1}}\left\lVert(f\circ u_{\textrm{enc}})^{{}^{\prime\prime}}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}.$ (13)
@@ A-A Proof of Theorem 1
$\stackrel{{\scriptstyle\text{}}}{{=}}\frac{2}{K}\sum^{K}_{k=1}\mathrm{E}_{\mathcal{F}}\left[h(\alpha_{k})^{2}\right]$
$\stackrel{{\scriptstyle\text{}}}{{\leqslant}}\frac{2}{K}\sum^{K}_{k=1}\mathrm{E}_{\mathcal{F}}\left[\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\right].$ (19)
@@
$\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\leqslant 2\left\lVert h\right\rVert_{{L}^{2}\left(\Omega\right)}\cdot\left\lVert h^{\prime}\right\rVert_{{L}^{2}\left(\Omega\right)}.$ (20)
Thus, we have:
$\mathrm{E}_{\mathcal{F}\sim F_{N}}\left[\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\right]$ $\leqslant 2\mathrm{E}_{\mathcal{F}}\left[\left\lVert h\right\rVert_{{L}^{2}\left(\Omega\right)}\left\lVert h^{\prime}\right\rVert_{{L}^{2}\left(\Omega\right)}\right]$
$\stackrel{{\scriptstyle\text{(a)}}}{{\leqslant}}2\mathrm{E}_{\mathcal{F}}\left[\left\lVert h\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\right]^{\frac{1}{2}}\cdot\mathrm{E}_{\mathcal{F}}\left[\left\lVert h^{\prime}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\right]^{\frac{1}{2}},$ (21)
where (a) follow from Cauchy-Schwartz inequality.
@@
Combining results in (22) and (23) with (A-A), we can conclude:
$\mathrm{E}_{\mathcal{F}}\left[\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\right]$ $\leqslant 2\sqrt{H_{0}H_{1}}\left\lVert(f\circ u_{\textrm{enc}})^{{}^{\prime\prime}}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\cdot\mathrm{E}_{\mathcal{F}}\left[L\right]^{\frac{1}{2}}\mathrm{E}_{\mathcal{F}}\left[L^{\frac{1}{2}}(1+\frac{L}{16})^{\frac{1}{2}}\right]^{\frac{1}{2}}.$ (26)
@@ A-B Proof of Theorem 2
$\mathrm{E}_{\mathcal{F}}\left[\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\right]$ $\stackrel{{\scriptstyle\text{}}}{{\leqslant}}\mathrm{E}_{\mathcal{F}}\left[\left(\Delta^{\mathcal{F}}_{\textrm{max}}\left(1+\frac{\Delta^{\mathcal{F}}_{\textrm{max}}}{\Delta^{\mathcal{F}}_{\textrm{min}}}\right)\right)^{2}\right]\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}$
$\stackrel{{\scriptstyle\text{}}}{{\leqslant}}\mathrm{E}_{\mathcal{F}}\left[\left(\Delta_{\textrm{max}}\right)^{2}\cdot\left(R_{\mathcal{F},N}+1\right)^{2}\cdot\left(1+B\left(R_{\mathcal{F},N}+1\right)\right)^{2}\right]\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}$
$\stackrel{{\scriptstyle\text{(a)}}}{{\leqslant}}\mathrm{E}_{\mathcal{F}}\left[\frac{J^{2}\left(R_{\mathcal{F},N}+1\right)^{2}}{N^{2}}\cdot\left(1+B\left(R_{\mathcal{F},N}+1\right)\right)^{2}\right]\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}$
$\stackrel{{\scriptstyle\text{}}}{{\leqslant}}2J^{2}B\cdot\mathrm{E}_{\mathcal{F}}\left[\frac{\left(R_{\mathcal{F},N}+1\right)^{4}}{N^{2}}\right]\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2},$ (37)
@@ A-D Proof of Corollary 2
$\mathrm{E}_{\mathcal{F}}\left[\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\right]$ $\leqslant\mathrm{E}_{\mathcal{F}}\left[\left(\Delta^{\mathcal{F}}_{\textrm{max}}\right)^{2}(1+\mu)^{2}\right]\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}$
$\stackrel{{\scriptstyle\text{(a)}}}{{\leqslant}}\mathrm{E}_{\mathcal{F}}\left[\left(\frac{J_{2}(R_{\mathcal{F},N}+1)}{N}\right)^{2}(1+\frac{((R_{\mathcal{F},N}+1)+1)((R_{\mathcal{F},N}+1)+3)\pi^{2}}{4})^{2}\right]\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}$
$\stackrel{{\scriptstyle\text{}}}{{\leqslant}}\mathrm{E}_{\mathcal{F}}\left[\tilde{C}\frac{(R_{\mathcal{F},N}+1)^{4}}{N^{2}}\right]\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2},$ (44)

### 24 Mar 2026 v3

@@ Abstract
Coded computing has demonstrated promising results in addressing straggler resiliency in distributed computing systems. However, most coded computing schemes are designed for exact computation, requiring the number of responding servers to exceed a certain recovery threshold. Additionally, these schemes are tailored for highly structured functions. Recently, new coded computing schemes for general computing functions, where exact computation is replaced with approximate computation, have emerged. In these schemes, the availability of additional results corresponds to more accurate estimation of computational tasks. This flexibility introduces new questions that need to be addressed. This paper addresses the practically important scenario in the context of general coded computing, where each server may become a straggler with a probability $p$, independently from others. We theoretically analyze the approximation error of two existing general coded computing schemes: Berrut Approximate Coded Computing (BACC) and Learning Theoretic Coded Computing (LeTCC). Under the probabilistic straggler configuration, we demonstrate that the average approximation error for BACC and LeTCC converge to zero with high probability with the rate of at least $\mathcal{O}(\log^{3}_{\nicefrac{{1}}{{p}}}(N)\cdot{N^{-3}})$ and $\mathcal{O}(\log^{4}_{\nicefrac{{1}}{{p}}}(N)\cdot{N^{-2}})$, respectively. This is perhaps surprising, as earlier results does not indicate a convergence when the number of stragglers scales with the total number of servers $N$. However, in this case, despite the average number of stragglers being $Np$, the independence of servers in becoming stragglers allows the approximation error to converge to zero. These theoretical results are validated through experiments on various computing functions, including deep neural networks.
@@ I Introduction
In this paper, we address this question and show that the above understanding is not correct. We theoretically analyze the approximation error in the presence of probabilistic stragglers for two general coded computing schemes of BACC and LeTCC. We demonstrate that the approximation errors of BACC and LeTCC schemes converge to zero at rates of at least $\mathcal{O}(\log^{3}_{1/p}(N)N^{-3})$ and $\mathcal{O}(\log^{4}_{1/p}(N)N^{-2})$, respectively, with high probability. This result demonstrates that the independence of servers in becoming stragglers indeed contributes positively to the convergence of the error. In the end, we validate our theoretical findings through experiments on various computing functions, including deep neural networks (see Fig. 1).
@@ III Problem Formulation
Following the notation in [6], let $\mathbf{\hat{f}}_{\bm{\alpha},\bm{\beta}}[\mathbf{u}_{\textrm{enc}},\mathbf{u}_{\textrm{dec}},\mathcal{F}](\cdot)$ denote the estimator function for the coded computing scheme. Let $\bm{\alpha}=[\alpha_{1},\dots,\alpha_{K}]^{T}$ and $\bm{\beta}=[\beta_{1},\dots,\beta_{N}]^{T}$. Furthermore, let $\mathcal{F}\sim F_{N}$ denote a random variable representing the set of non-straggler servers, where $F_{N}$ is a distribution over all subsets of $N$ servers. Consequently, the performance of a coded computing estimator $\mathbf{\hat{f}}(\mathbf{x}):=\mathbf{\hat{f}}_{\bm{\alpha},\bm{\beta}}[\mathbf{u}_{\textrm{enc}},\mathbf{u}_{\textrm{dec}},\mathcal{F}](\cdot)$, applied to the data points $\{\mathbf{x_{k}}\}_{k=1}^{K}$, can be evaluated using the following average approximation error:
@@ IV Main Results
$\mathcal{L}_{\mathcal{F}}(\hat{f})\leqslant C_{l}\frac{\left(\log_{\frac{1}{p}}(qN)+\sqrt{\frac{1}{\delta}}\right)^{3}}{N^{3}}\cdot\left\lVert(f\circ u_{\textrm{enc}})^{\prime\prime}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}+\frac{2\nu^{2}}{K}\sum_{k=1}^{K}\left(f(u_{\textrm{enc}}(\alpha_{k}))-f(x_{k})\right)^{2}.$ (7)
@@
$\mathcal{L}_{\mathcal{F}}(\hat{f}){\leqslant}C_{b}\cdot B\frac{\left(\log_{\frac{1}{p}}(qN)+\sqrt{\frac{1}{\delta}}\right)^{4}}{N^{2}}\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}.$ (8)
@@
By applying Theorems 1 and 2, in the probabilistic straggler setting, the average approximation error converges to zero with high probability. The convergence rates are at least $\mathcal{O}(\frac{\log^{3}_{1/p}(N)}{N^{3}})$ for LeTCC and at least $\mathcal{O}(\frac{\log^{4}_{1/p}(N)}{N^{2}})$ for BACC.
@@ V Proof Sketch
$\mathcal{L}_{\mathcal{F}}(\hat{f})\leqslant\underbrace{\frac{2}{K}\sum^{K}_{k=1}\left(u_{\textrm{dec}}(\alpha_{k})-f(u_{\textrm{enc}}(\alpha_{k}))\right)^{2}}_{\mathcal{L}_{\textrm{dec}}(\hat{f})}+\underbrace{\frac{2}{K}\sum^{K}_{k=1}\left(f(u_{\textrm{enc}}(\alpha_{k}))-f(\mathbf{x}_{k})\right)^{2}}_{\mathcal{L}_{\textrm{enc}}(\hat{f})}$. (9)
@@
$\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\leqslant 2\left(\left\lVert h\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\cdot\left\lVert h^{\prime}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\right)^{\frac{1}{2}}.$
@@
$\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\leqslant C\cdot D\mathrm{E}_{\mathcal{F}}\left[(\Delta^{\mathcal{F}}_{\textrm{max}})^{4}\right]^{\frac{1}{2}}\cdot\sqrt{D}\mathrm{E}_{\mathcal{F}}\left[(\Delta^{\mathcal{F}}_{\textrm{max}})^{2}\right]^{\frac{1}{2}}\cdot\sqrt{H_{0}H_{1}}\left\lVert(f\circ u_{\textrm{enc}})^{{}^{\prime\prime}}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}.$ (13)
@@ A-A Proof of Theorem 1
$\stackrel{{\scriptstyle\text{}}}{{=}}\frac{2}{K}\sum^{K}_{k=1}h(\alpha_{k})^{2}$
$\stackrel{{\scriptstyle\text{}}}{{\leqslant}}\frac{2}{K}\sum^{K}_{k=1}\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}.$ (19)
@@
$\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\leqslant 2\left(\left\lVert h\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\cdot\left\lVert h^{\prime}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\right)^{\frac{1}{2}}.$ (20)
@@
Combining results in (21) and (22) with (20), we can conclude:
$\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}$ $\leqslant 2\sqrt{H_{0}H_{1}}\left\lVert(f\circ u_{\textrm{enc}})^{{}^{\prime\prime}}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\cdot L^{\frac{1}{2}}\left[L^{\frac{1}{2}}(1+\frac{L}{16})^{\frac{1}{2}}\right]^{\frac{1}{2}}.$ (25)
@@ A-B Proof of Theorem 2
$\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}$ $\stackrel{{\scriptstyle\text{}}}{{\leqslant}}\left(\Delta^{\mathcal{F}}_{\textrm{max}}\left(1+\frac{\Delta^{\mathcal{F}}_{\textrm{max}}}{\Delta^{\mathcal{F}}_{\textrm{min}}}\right)\right)^{2}\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}$
$\stackrel{{\scriptstyle\text{}}}{{\leqslant}}\left(\Delta_{\textrm{max}}\right)^{2}\cdot\left(R_{\mathcal{F},N}+1\right)^{2}\cdot\left(1+B\left(R_{\mathcal{F},N}+1\right)\right)^{2}\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}$
$\stackrel{{\scriptstyle\text{(a)}}}{{\leqslant}}\frac{J^{2}\left(R_{\mathcal{F},N}+1\right)^{2}}{N^{2}}\cdot\left(1+B\left(R_{\mathcal{F},N}+1\right)\right)^{2}\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}$
$\stackrel{{\scriptstyle\text{}}}{{\leqslant}}2J^{2}B\cdot\frac{\left(R_{\mathcal{F},N}+1\right)^{4}}{N^{2}}\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2},$ (36)
@@ A-D Proof of Corollary 2
$\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}$ $\leqslant\left(\Delta^{\mathcal{F}}_{\textrm{max}}\right)^{2}(1+\mu)^{2}\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}$
$\stackrel{{\scriptstyle\text{(a)}}}{{\leqslant}}\left(\frac{J_{2}(R_{\mathcal{F},N}+1)}{N}\right)^{2}(1+\frac{((R_{\mathcal{F},N}+1)+1)((R_{\mathcal{F},N}+1)+3)\pi^{2}}{4})^{2}\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}$
$\stackrel{{\scriptstyle\text{}}}{{\leqslant}}\tilde{C}\frac{(R_{\mathcal{F},N}+1)^{4}}{N^{2}}\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2},$ (43)